8. The mean number of patient visits at a clinic is 5.6 per week. to find the probability that Use the Poisson probability diplease answer all parts
Answer #1


#8: mean number (=5.6; pla)= (@s.65 €3) 2. Ca pir=2) = (sobre -10.05798) (5) Pla>5) = l-pra<s) = 1- $piazo) +pIm=) +P/1=2)+P